OpenStudy (anonymous):

a+a(1,15) + a (1,15)^2 + a (1,15)^3+ ..........+a (1,15)^11 = 2900 How does I find a ?

OpenStudy (anonymous):

What does (1,15) refers to.

OpenStudy (anonymous):

??

OpenStudy (anonymous):

??

OpenStudy (anonymous):

I dont know :( I just have to fint the value of a, i dont understand how

OpenStudy (anonymous):

something is missing in the question...

OpenStudy (anonymous):

Nope this is the text in the task...

OpenStudy (anonymous):

wat is the topic... ?? this is related to what??

OpenStudy (anonymous):

Decide a when \[a +a \times 1,15 + a \times(1,15)^{2} + a \times (1,15)^{3} etc........ + a (1,15)^{11} = 2900\]

OpenStudy (anonymous):

geometric series